Physics – Geometrical Optics-
A man is looking down a 9 m deep tank filled with a transparent liquid of refractive index √3 . The line of sight of the man makes an angle of 30° with the horizontal. What is the apparent depth of the tank as seen by the man? 
 

Grade:12

1 Answers

Arun
5 years ago
Real Depth of the tank = 9 m.
Refractive Index of the liquid(μ) = √3
∵ μ = Real Depth/Apparent Depth
∴ Apparent Depth = Real Depth/μ
∴ Apparent Depth = 9/√3
∴ Apparent Depth = 9/√3 × √3/√3
⇒ A.D. = 3√3 m.
⇒ A.D. = 3 × 1.732
⇒ A.D. = 5.196 m.

Thus, the depth of the tank as seen by the man is 5.196 m.
___________________

Also,angle made by the line of sight with the Horizontal(or angle of Incidence) = 30°
Now, μ =Sin i/Sin r
⇒ √3 = Sin 30°/Sin r
⇒ Sin r = 1/2 ÷ √3
⇒ Sin r = √3/2
⇒ Sin r = Sin 60°
On Comparing,
  r = 60°
Thus, angle of Refraction is 60°.
 
 
Regards
Aru. (askIITians forum expert)